Agenda/Minutes on 2020/12/11
CRY+VIS Interface Meeting
Participants: Takahashi, Miyo, Ikeda, Aso, Ushiba, Yano, Sato, Hirata
- Schedule: Connection of ITMX after 1/12. Experiments of ITMY are delay. The schedule will be changed.
- Flow chart: Two-step commissioning in vacuum were suggested.
- Acceptance check items: Making the document for cryo-payload.
- Jig for FC painting: Asking quotation for the application to both sides.
PF actuator: Measurement of coefficients --> calculation of the noise (by the end of Jan.).
Participants: Takahashi, Miyo, Aso, Yano, Terrence, Fabian, Hirata, Sato, Tanaka
- Modified three ACCs for ITMX. The capacitance sensors were replaced to the LVDTs.
Week of the 7th of December
- We connected some SR2 and SRM cables which had been disconnected as preparation for the clean booth cleaning.
- We BS IP immobilizing jig and LVDT-actuator units.
- Some holes were out of place due to manufacturing imperfections but we were able to fix the component using only one screw.
- Then we found another hole was also out of place. Hirata-san and Yano-san made additional holes and it worked very well.
- After immobilizing the IP we removed the LVDTs and diassemebled them.
- The LVDTs had oil, rust, dust, ink, small pieces of polystyrene foam, etc.
- We wrote a rough procedure to use it as a guide to assemble them back, including cabling diagram.
- Some of the think wires broke when removing the pins from the connectors, but Hirata-san crimpped new pins successfully.
- We measured the resistances of the coils and determined the polarity of magnet arrangement.
- The yokes were given to Takahashi-san who sent them to the company. We'll have them back by the 23rd of December.
- On Tuesday and Wednesday Hirata-san and Yano-san will go to Toyama University to wash the components.
- On Monday morning Yano-san will compile a bill of screws based on our notes.
- The BS F0 stepper motor was assembled but not installed. We'll do it after recovering the IP LVDTs.
- Two BS IP horizontal stepper motor limit switches were assembled.
- One of them was put in place just for us to check iy fitted in place. It looked alright.
- The assembly does not have cable clamps for the cables, but we thought of a plan. We only require to but additional screws.
- In SR3 IP we wanted to replace a heavy disc in the F0 keystone assembly. However, the holes in the new piece have countersinks and current piece uses socket head screws. We will install it after buying new suitable screws.
- Yano-san made some holes bigger in some ballast masses (from M5 to M6).
- The temperature in BS area is currently 23.8 degrees. We want around 23 degrees. Hopefully next week we will be able to adjust the ballast mass in BS BF.
- I began writing the new procedure to remove ballast mass from SR3 Im inside of the chamber.
- No need to design new ballast masses or any other piece.
- I aim to finish it shortly and I'll ask Hirata-san to check it.
- I should be able to tell you how long it will take to do this some time next week.
- OMMT1 and OMMT2 were recovered.
- Acceptance test is necessary.
- Results of prototype test for the combiner were shown. Signal from the secondary coil was too small (-20dB). So it was difficult to confirm effects due to the filter circuit.
- The circuit will be sent to Kamioka.
- Type-B remedy work is on going.
- Modification of ACCs is on going.
- Procedure documents.
- Criteria in the acceptance test.
Hirata-san got metal splinters in two fingers in the workshop inside the tunnel. They were removed more or less easily. Suitable gloves should be provided.
The week from 12/14
- Takahashi: 12/14-18
- Hirata: 12/14-18
- Yano: 12/14-18
(F2F Meeting in 12/17-18)
The week from 12/21
- Takahashi: 12/21-24
- Hirata: 12/21-24
- Yano: 12/21-24
- On 2020/12/25(Fri)